North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog 10th and 11th November 2020

10th and 11th November 2020

Ian reports a mild dull morning, but brightening up just before lunchtime, greeted Bill Dobson, Nigel Bill and Ian himself at Grosmont on Tuesday.

Nigel and Ian continued cleaning brake gear and spring components. Bill spent some time down at the MPD with Nick Simpson, arranging for the two new stepped studs required for the boiler backplate which hold the tea tray in position. While Nigel was cleaning springs, spring hangers and brake tables, Ian continued to clean the remaining six spring hanger bolts in the washer in the repair shed. However, the nut on the the left driving back spring hanger would not come free: it was stuck solid. After several attempts with a large spanner with an extension on its shaft, it still would not move. So the next couple of hours were spent in Deviation Shed trying to remove it with some heat. It was then soon found that the oxygen bottle was empty (the failing equipment curse strikes again!) so the bottle was replaced. This in itself took quite some time as the full ones are in a compound near the boiler shop. After setting up the burning gear with the new oxy bottle it was time for lunch, which was had in Deviation shed with Zack the dog.

After lunch, Nigel did the monthly Defibrillator checks (all OK) and Bill measured up the boiler so he can order the insulation. He also was in conversation with Barry Neesom regarding the safety valves which are almost ready. Ian and Nigel heated the nut on the spring hanger bolt in Deviation Shed and eventually, after about an hour, it finally came off. After it cooled down, it was cleaned, but, as it was still tight on the threads, Bill took it to Barry Neesham who gave a thump on the sides of the nut while screwed on the bolt and this allowed it to turn freely. While Ian cleaned the last two spring hanger bolts, Bill cleaned the paint from inside frames around the horn guides ready for an NDT inspection. Bill did this and Ian took the necessary photos which have been sent to Barney for the Q6 file. No cracks were detected. Sean Bowler had been working on his own shimming the horn guides on Monday and Tuesday, and his next job will be to measure the new distance between the horn guides against each axle box.

Wednesday was a rather dull cool morning but brighter later on, for the rostered Ed Bolam, Nigel Bill, Nigel Hall, Bill Dobson and Ian Pearson. The day was mainly spent cleaning and painting. Last Friday, Nigel Hall gave the boiler barrel another coat of silver paint, and today the fire box area got the same treatment to finish the job. Ed, Ian and Nigel Bill spent all day cleaning the wheel sets and axle box top plates, in the process and removing oil and water from under the top plates along with the old trimmings which will have to be replaced. Bill investigated the state of the underkeep pads. After dismantling the trailing wheel underkeep, the pad was found to be in very poor condition and, after discussing the issue with Barney, it was decided to change all 8 pads. Bill will be putting an order in for these at Armstrong Oilers shortly. Sean Bowler and Dave Brown have completed the shimming of the horn guides, and have fitted new nuts to the horn stays. Now they are all the same size nuts !

Two minutes silence was held at 11am in Deviation Shed and the MPD on this Remembrance Day.